Assume that a sample is used to estimate a population proportion μμ. Find the margin of error M.E. that corresponds to a sample of size 45 with a mean of 71.2 and a standard deviation of 8.9 at a confidence level of 99%.

Report ME accurate to one decimal place because the sample statistics are presented with this accuracy.
M.E. =

Answer should be obtained without any preliminary rounding. However, the critical value may be rounded to 3 decimal places.
Answer #1

given.

\bar{X}=sample mean=71.2

s=sample  standard deviation is= 8.9

n=45

M.E. =margin of error

ME = ta/2, (n − 1)*

ta/2, (n − 1) = 0.01/2,44 = 2.6922

value of t is obtain from t-dist with corresponding degree of freedom 44

8.9 ME = 2.6922 * 145

=3.572

ME=3.572

#